How far will a jumper go if he accelerates to 11.0 m/sec and jumps at an angle of 30
1
Expert's answer
2020-12-14T07:26:34-0500

Horizontal components

"Horizontal\ components
\\
V_{ix} = V_{i}\times \cos \Theta \\

V_{ix} = 11.0 m/s \times \cos \ 30\\
V_{ix} = 9.5 m/s 

\\
Vertical\ components

\\
V_{iy} = V_{i}\times \sin \Theta \\

V_{iy} = 11.0 m/s \times \sin \ 30\\
\\
V_{iy} = 6 m/s \\
To\ find\ time \ t \\
V_{fy} = V _{iy}+ a_{y}t\\
Subtituting \ the \ values\ in\ the\ above\ equation\\
-6m/s = 6m/s +(-9.8m/s^2)\\

Solving\ the\ equation\ we\ get\ time\ (t)\\
we\ get \\
t=1.22s\\
To\ get\ how\ far\ will\ the\ jumper\ go (X)\\
\\

X=V_{ix}t +\frac{1}{2}a_{x}t^{2}\\
X=9.5m/s\times1.22s+\frac{1}{2}(0m/s^2)\times(1.22)^2\\
X=11.59m"